BexleyBrooks:

Need some more help on Math. 1. 5+4(n+9)=-3 what is N? 2. 5(k-2)-8k=-34 what is K? 3. 2=11+3(m+3) what is M? 4. -2(p-5)+7p=-5 what is P? 5. 4a-2a(a+9)=6 what is A? 6. 7-4(d-3)=23 what is D? 7. 8x-11(x-2)=-8 what is X? NO ARGUING ON THIS ONE! I WILL NOT TOLERATE IT!

BexleyBrooks:

@hotpocket14 this is a new post since everyone kept arguing on the other one

xXChudLord420Xx:

*argugigng*

luhivqqcherry:

1. 5+4(n+9)=-3 what is N? n=11

HotPockets14:

1).N = -11 2).K = 8 3). M = -6 4). P = 5/9 5). A = -7+37/2 6). D = -1 7). X = 10

BexleyBrooks:

@hotpockets14 wrote:
1).N = -11 2).K = 8 3). M = -6 4). P = 5/9 5). A = -7+37/2 6). D = -1 7). X = 10
I don't get how you got -7+37/2 for number 5. Could you maybe explain it to me a little more?